OpenStudy (anonymous):

f(x)=-3/2cos(3x+4pi)+5 Find The Period, Range, Phase Shify, and Vertical Shift.

OpenStudy (anonymous):

@roadjester

OpenStudy (roadjester):

\(\Large f(x) = -\dfrac {3} 2 cos(3x+4\pi)\)+5 correct?

OpenStudy (anonymous):

Yes sir

OpenStudy (roadjester):

ok so standard form is a*cos(bx+c)+d

OpenStudy (anonymous):

Isn't already in that form?..

OpenStudy (roadjester):

exactly

OpenStudy (roadjester):

now, once it's in that form, you're given all the information you were asked to look for

OpenStudy (roadjester):

let's start with the easiest two, amplitude and vertical shift

OpenStudy (roadjester):

amplitude of a sin/cos function is simply "a" and the vertical shift is simply "d"

OpenStudy (anonymous):

Amplitude is -3/2

OpenStudy (roadjester):

make sense

OpenStudy (anonymous):

Si senor

OpenStudy (roadjester):

whoops, ok my mistake

OpenStudy (anonymous):

Go Get Your Dinner

OpenStudy (roadjester):

amplitude is |a|

OpenStudy (roadjester):

so in this case, the graph will go up 1.5 units and down 1.5 units

OpenStudy (roadjester):

the negative sign means it's been flipped upside -down

OpenStudy (anonymous):

OOOhhh, Got You

OpenStudy (anonymous):

The vertical shift is 5?

OpenStudy (roadjester):

ok the vertical shift itself depends only on the |d|, but the direction it is shifted vertically depends on the sign

OpenStudy (anonymous):

Ummm...

OpenStudy (roadjester):

think of it this way, when you have y=mx+b

OpenStudy (roadjester):

the b is your y intercept and it is also your vertical shift; make sense?

OpenStudy (anonymous):

only if i could understand

OpenStudy (roadjester):

umm ok have you ever seen a sine or cosine wave before?

OpenStudy (anonymous):

so is it 4

OpenStudy (roadjester):

 That's the shape of your standard cosine wave